# finite free probability and $S$ transforms of Jacobi processes

### v1

@@ 5.2. Finite free probability and finite differences of the finite free $S$ transforms
The goal of this section is to propose another derivation of the PDE satisfied by the free $S$ transform of the Jacobi process, see Theorem 3.1.1, by using finite free probability and the finite free $S$-transform. It requires a technical result regarding the asymptotics of the discrete derivatives of the finite free $S$ transform, which is proven in the following paragraphs. The reader will find all relevant notions and definitions of objects used in this section in the series of papers [AP18, AP18, MSS22]. For brevity, only the basic definitions are recalled here.
@@ 5.2.1. The finite free $S$ and $T$ transforms
Let $p$ be a monic polynomial with degree $p$,
$p=\sum_{k=0}^{d}(-1)^{k}e^{(d)}_{k}(p)x^{n-k}.$
Call $r$ the multiplicity of 0 in $p$ and assume that $p$ has only positive real roots. Then the finite free $S$ transform of $p$, denoted hereafter by $S^{(d)}_{p}$ is the function on the set of points $\{-k/d,k\in\{1,d-r\}\}$ :
$\mathcal{S}_{p}^{(d)}\left(-\frac{k}{d}\right):=\frac{d-k+1}{k}\frac{e^{(d)}_{k-1}(p)}{e^{(d)}_{k}(p)}.$
It will also be convenient to introduce the finite free $T$ transform of $p$ : it is the piecewise right-continuous function over $(0,1)$ defined by
$T_{p}^{(d)}(z)=\begin{cases}0,&z\in(0,\frac{r}{d})\\ \frac{d-k+1}{k}\frac{e^{(d)}_{d-k+1}(p)}{e^{(d)}_{d-k}(p)},&z\in[\frac{k-1}{d},\frac{k}{d}),\quad k=r+1\ldots d\end{cases}$
@@ 5.2.2. Convergence of the finite differences
We let $\nabla^{(d)}$ be the operator of finite right-differentiation with step $\frac{1}{d}$ acting on functions of $\mathbb{R}$:
$\nabla^{(d)}\colon\mathbb{R}^{\mathbb{R}}\to\mathbb{R}^{\mathbb{R}},\quad\nabla^{(d)}g(v):=d(g(v+1/d)-g(v)),\quad v\in\mathbb{R}.$
Before stating the next Theorem, recall that if $(p_{d})_{d\geq 1}$ is a sequence of polynomials with increasing degree $d$, we denote by $\mu\llbracket p_{d}\rrbracket$ the counting measure of its roots :
$\mu\llbracket p_{d}\rrbracket:=\frac{1}{d}\sum_{z\,:\,P(z)=0}\delta_{z}$
[…]
@@
Let $(p_{d})_{d\geq 1}$ be a converging sequence of monic polynomials with positive roots and increasing degrees $d$. Let $\mu$ be the limiting measure :
$\mu\llbracket p_{d}\rrbracket\underset{d,+\infty}{\longrightarrow}\mu.$
Assume that $\mu\neq\delta_{0}$. Let $v\in(\mu(\{0\}),1)$. Then, as $d$ goes to infinity,
$[\nabla^{(d)}](T^{(d)}_{p_{d}})(v)=\partial_{v}T_{\mu}(v)+o_{d}(1)$
locally uniformly on $v$.
@@
We will prove the result for the left-continuous step-function $\mathcal{S}^{(d)}_{p_{d}}$ interpolating over the interval $(0,1)$ the finite free $S$-transform of the polynomials $p_{d}$:
$\mathcal{S}^{(d)}_{p_{d}}\colon(0,1)\to\mathbb{R},\quad\mathcal{S}^{(d)}_{p_{d}}(v)=S^{(d)}_{p_{d}}(-\frac{\lceil{dv}\rceil}{d}),\,v\in(0,1-r_{d}/d),$
where we recall that $\lceil{\cdot}\rceil$ is the ceiling function and $r_{d}$ is the multiplicity of 0 in $p_{d}$. It will then be straightforward to transform the statement about the $S$-transform to the same statement on $T$ transform given that, for any $v\in(0,1)$:
$T_{p}^{(d)}(1-v)=\mathcal{S}^{(d)}_{p_{d}}(-v)^{-1}.$
We closely follow the line of arguments presented in [Ari+24] to prove Propositions 7.2, 7.3 and 7.8.
Recall that ${\rm Dil_{v}}\mu$ is the dilation by $v$ of the measure $\mu$: the $\ell^{th}$ moment of ${\rm Dil_{v}}\mu$ is $v^{\ell}$ times the $\ell$ moment of $\mu$. The fractional additive convolution power $\mu^{\boxplus\frac{1}{v}}$ of $\mu$ have its $\ell^{th}$ free cumulant equal to $\frac{1}{\ell}$ times the $\ell^{th}$ free cumulant of $\mu$.
@@
Case 1: We make the additional assumption that there exists $\varepsilon>0$ and $\eta>0$ such that all the roots of the polynomials $p_{d}$ are contained in $\varepsilon,\eta)$.
@@
We begin by relating the finite free $S$-transform to the Cauchy transform of the counting measure of the roots of derivatives of $p_{d}$, (see [Ari+24] Lemma 6.11 2). Let $v\in(0,1)$.
$d(\mathcal{S}^{(d)}_{p_{d}}(v+\frac{1}{d})-\mathcal{S}_{p_{d}}^{(d)}(v))$ $=d({S}^{(d)}_{p_{d}}(-\frac{\lceil{dv}\rceil+1}{d})-{S}^{(d)}_{p_{d}}(-\frac{\lceil{dv}\rceil}{d}))$
$=d(G_{\mu\llbracket\partial^{\lceil{dv}\rceil|}p_{d}\rrbracket}(0)-G_{\mu\llbracket\partial^{\lceil{dv}\rceil+1|d}p_{d}\rrbracket}(0))$
$=G_{\nu^{(d)}_{p_{d}}(v)}(0),$
where $\nu^{(d)}_{p_{d}}(v)$ is the signed measure (with atomic support) defined by
$\nu^{(d)}_{p_{d}}(v)=\frac{d}{\lceil{dv}\rceil}\sum_{\lambda\in Z(\partial^{\lceil{dv}\rceil|d}p_{d})}\delta_{\lambda}-\frac{d}{\lceil{dv}\rceil+1}\sum_{\lambda\in Z(\partial^{\lceil{dv}\rceil+1|d}p_{d})}\delta_{\lambda}.$
@@
Observe that this measure has total variation equal to $2d$, it is therefore not uniformly bounded. Nevertheless, we will study the convergence of its Cauchy transform by using the moment method.
By definition, the $\ell^{th}$ moment $m_{\ell}(\nu_{p_{d}}^{(d)}(v))$ of $\nu_{p_{d}}^{(d)}(v)$ is
$m_{\ell}(\nu_{p_{d}}^{(d)}(v))$ $=d(m_{\ell}(\mu\llbracket\partial^{\lceil{dv}\rceil|d}p_{d}\rrbracket)-m_{\ell}(\mu\llbracket\partial^{\lceil{dv}\rceil+1|d}p_{d}\rrbracket))$
$=d(m^{(d)}_{\ell}(\partial^{\lceil{dv}\rceil|d}p_{d})-m^{(d)}_{\ell}(\partial^{\lceil{dv}\rceil+1|d}p_{d}))$ (18)
@@
Recall the following asymptotic expansion between the moments $m^{(d)}_{\ell}(p_{d}),\,\ell\geq 1$ and the finite free cumulants $\kappa^{(d)}_{\ell}(p_{d}),\ell\geq 1$ of $p_{d}$ (see [AP18]):
$m_{\ell}^{(d)}(p)=\sum_{\pi\in NC(\ell)}\kappa^{(d)}_{\pi}(p)+\frac{2\ell}{d}\sum_{\begin{subarray}{c}r+s=\ell\\ \pi\in S_{NC}(r,s)\end{subarray}}r^{-1}s^{-1}\kappa^{(d)}_{\pi}(p)+O(\frac{1}{d^{2}})$ (19)
@@
and the following formula relating the finite free cumulant of a polynomial $p$ and its derivative $\partial^{n|d}p$ (see [Ari+24]) :
$\kappa^{(k)}_{\ell}(\partial^{k|d}(p))=(\frac{k}{d})^{\ell-1}\kappa^{(d)}_{\ell}(p),\quad\ell\leq k.$ (20)
Inserting (19) in (5.2.2) and using further equation (20), we obtain:
$m_{\ell}(\nu^{(d)}_{p_{d}}(v))$ $=d\bigg(\sum_{\pi\in NC(\ell)}\bigg[(\frac{\lceil{vd}\rceil+1}{d})^{\ell-|\pi|}\kappa^{(d)}_{\pi_{d}}(p_{d})-(\frac{\lceil{vd}\rceil}{d})^{\ell-|\pi|}\kappa_{\pi}^{(d)}(p_{d})\bigg]$
$\hskip 56.9055pt+2\frac{\ell}{\lceil{vd}\rceil+1}\sum_{\begin{subarray}{c}r+s=\ell\\ \pi\in S_{NC}(r,s)\end{subarray}}r^{-1}s^{-1}(\frac{\lceil{vd}\rceil+1}{d})^{\ell-|\pi|}\kappa^{(d)}_{\pi}(p_{d})$
@@
$\hskip 142.26378pt-2\frac{\ell}{\lceil{vd}\rceil}\sum_{\begin{subarray}{c}r+s=\ell\\ \pi\in S_{NC}(r,s)\end{subarray}}r^{-1}s^{-1}(\frac{\lceil{vd}\rceil}{d})^{\ell-|\pi|}\kappa_{\pi}^{(d)}(p_{d})\bigg)$
$+O(1/d)$
@@
For the first term in the right-hand side of the equation above, with $\kappa_{\pi}(\mu),\pi\in{\rm NC}(\ell)$ the partitioned free cumulants of $\mu$:
$d\sum_{\pi\in NC(\ell)}\bigg[(\frac{\lceil{vd}\rceil+1}{d})^{\ell-|\pi|}$ $\kappa^{(d)}_{\pi}(p)-(\frac{\lceil{vd}\rceil}{d})^{\ell-|\pi|}\kappa_{\pi}^{(d)}(p)\bigg]$
$=d\sum_{\pi\in NC(\ell)}\bigg[(\frac{\lceil{vd}\rceil}{d})^{\ell-|\pi|}((1+\frac{1}{\lceil{vd}\rceil})^{\ell-|\pi|}-1)\kappa^{(d)}_{\pi}(p_{d})\bigg]$
$=\sum_{\begin{subarray}{c}\pi\in NC(\ell)\\ \pi\neq\hat{0}_{\ell}\end{subarray}}\bigg[(\ell-|\pi|)v^{\ell-|\pi|-1}\kappa_{\pi}(\mu)\bigg]+O(1/d)$
@@
The sum of the second term and the third term are easily seen to contribute to a factor $O(\frac{1}{d})$. Hence, we obtain for the $\ell^{th}$ moment of $\nu_{p_{d}}^{(d)}(v)$ the following asymptotic expansion:
$m_{\ell}(\nu_{p_{d}}^{(d)}(v))$ $=\sum_{\begin{subarray}{c}\pi\in NC(\ell)\\ \pi\neq\hat{0}_{\ell}\end{subarray}}\bigg[(\ell-|\pi|)v^{\ell-|\pi|-1}\kappa_{\pi}(\mu)\bigg]+O({1}/{d})$
$=\partial_{v}\sum_{\pi\in NC(\ell)}v^{\ell-|\pi|}\kappa_{\pi}(\mu)+O(1/d)$
$=\partial_{v}m_{\ell}({\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}})+O(1/d)$
$=m_{\ell}(\partial_{v}{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}})+O(1/d)$
@@
Let us argue now that the Cauchy transform of $\nu_{p_{d}}^{(d)}(v)$ converges to the Cauchy transform of $\partial_{v}{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}}$ (analytical in the complementary of an interval $\varepsilon(v),\eta(v)]$ with $\varepsilon(v)>0$ (resp. $\eta(v)$) as a consequence of Lemma 5.6 and the proof of Proposition 5.5 in [Ari+24]; in the sequel $\varepsilon$ (resp. $\eta$) will denote the minimum (resp. maximum) between $\varepsilon(v)$ and $\varepsilon>0$) (resp. between $\eta$ and $\eta(v)$). We observe first that:
$G_{\nu_{p}(v)}=\{G_{\nu^{(d)}_{p_{d}}(v)},\,d\geq 1\}$
@@
is a normal family of holomorphic functions on the domain $\Omega=[\varepsilon,1]^{c}\subset\hat{\mathbb{C}}$$\mathbb{C}\cup\{$$\infty\}$, by applying Montel’s criterion (Fundamental Normality Test). Let $G\colon\Omega\to\hat{\mathbb{C}}$ be an accumulation point of $G_{\nu_{p}(v)}$. First, $G$ is not identically equal to $\infty$ since $G_{\nu^{(d)}_{p_{d}}(v)}(\infty)=0$ for all $d\geq 1$. Since
$\frac{1}{\ell+1!}\frac{d^{\ell+1}}{dz^{\ell+1}}\biggr|_{z=0}G_{\nu^{(d)}_{p_{d}}(v)}(1/z)=m_{\ell}(\nu_{p_{d}^{(d)}})\rightarrow m_{\ell}(\partial_{v}{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}}),$

### v3

@@ 5.2. Finite free probability and finite differences of the finite free $S$ transforms
The goal of this section is to propose another derivation of the PDE satisfied by the free $S$ transform of the Jacobi process, see Theorem 3.1.1, by using finite free probability and the finite free $S$ transform. This derivation paves the way to study more general particles systems, beyond the compact case; such as the frozen Heckman-Opdam process of type BC [MR4406927]. It requires a technical result regarding the asymptotics of the discrete derivatives of the finite free $S$ transform, which is proven in the following paragraphs. The reader will find all relevant definitions used in this section in the series of papers [arizmendi2018cumulants, arizmendi2018cumulants, MR4408504]. For brevity, only the basic ones are recalled below.
@@ 5.2.1. The finite free $S$ and $T$ transforms
Let $p$ be a monic polynomial with degree $m$:
$p_{m}(x)=\sum_{k=0}^{m}(-1)^{k}e^{(m)}_{k}(p)x^{m-k},\quad e^{(m)}_{0}(p)=1.$
Let $r$ the multiplicity of 0 in $p_{m}$ and assume that $p_{m}$ has only positive real roots. Then the finite free $S$ transform of $p_{m}$, denoted hereafter by $S^{(m)}_{p_{m}}$, is the function defined on the set of points $\{-k/m,k\in\{1,m-r\}\}$ defined by [arizmendi2024s]:
${S}_{p_{m}}^{(m)}\left(-\frac{k}{m}\right):=\frac{m-k+1}{k}\frac{e^{(m)}_{k-1}(p)}{e^{(m)}_{k}(p)},\quad k\in\{1,m-r\}.$
It will also be convenient to introduce the finite free $T$ transform of $p_{m}$ : it is the piecewise right-continuous function defined on the open interval $(0,1)$ by
$T_{p_{m}}^{(m)}(v)=\begin{cases}0,&v\in(0,\frac{r}{m})\\ \displaystyle\frac{m-k+1}{k}\frac{e^{(m)}_{m-k+1}(p)}{e^{(m)}_{m-k}(p)},&v\in[\frac{k-1}{m},\frac{k}{m}),\quad k=r+1,\ldots,m.\end{cases}$
Note in passing that while the finite free S transform is defined on a finite set of real numbers, the finite free T transform is defined in the open interval $(0,1)$. This difference stems from the fact that the definition of the latter is motivated by the finite free analogue of the central limit Theorem for the free multiplicative convolution (see Remark 6.4 in [arizmendi2024s]). Nonetheless, for any $k\in\{1,\dots,m-r\},$
$T_{p_{m}}^{(m)}\left(\frac{m-k}{m}\right)=\frac{1}{\displaystyle{S}_{p_{m}}^{(m)}\left(-\frac{k}{m}\right)}.$ (19)
@@ 5.2.2. Convergence of the finite differences
We let $\nabla^{(m)}$ be the operator of finite right-differentiation with step $1/m$ acting on functions of $\mathbb{R}$:
$\nabla^{(m)}\colon\mathbb{R}^{\mathbb{R}}\to\mathbb{R}^{\mathbb{R}},\quad\nabla^{(m)}g(v):=m\left(g\left(v+\frac{1}{m}\right)-g(v)\right),\quad v\in\mathbb{R}.$
Likewise, the discrete left-derivative of $g$ is
$\tilde{\nabla}^{(m)}\colon\mathbb{R}^{\mathbb{R}}\to\mathbb{R}^{\mathbb{R}},\quad\tilde{\nabla}^{(m)}g(v)=m\left(g(v)-g\left(v-\frac{1}{m}\right)\right),\quad v\in\mathbb{R}$
Most of the following statements will be formulated by using the right-derivative operator $\nabla^{(m)}$, although all have their counterparts valid for the left derivative operator $\tilde{\nabla}$.
Before stating the next Theorem, recall that if $(p_{m})_{m\geq 1}$ is a sequence of polynomials with increasing degree $m$, we denote by $\mu\llbracket p_{m}\rrbracket$ the counting measure of its roots :
$\mu\llbracket p_{m}\rrbracket:=\frac{1}{m}\sum_{z\,:\,P(z)=0}\delta_{z}.$
If $\mu\llbracket p_{m}\rrbracket$ converges weakly to some probability measure $\mu$, we say that $(p_{m})_{m\geq 1}$ is a converging sequence. From [arizmendi2024s] (see Theorem 1.1. there), if for any $m\geq 1$, $p_{m}$ has only positive roots and if $\mu\neq\delta_{0}$, then $\mu\llbracket p_{m}\rrbracket$ converges to $\mu$ if and only if $(S^{(m)}_{p_{m}})_{m}$ converges to the free $S$-transform $S_{\mu}$ of $\mu$ in the following sense: for any $v\in(0,1-\mu\{0\})$ and any sequence $(k_{m})_{m\geq 1}$ such that
$1\leq k_{m}\leq m,\quad\lim_{m\rightarrow+\infty}\frac{k_{m}}{m}=t,$
one has:
$\lim_{m\rightarrow+\infty}{S}_{p_{m}}^{(m)}\left(-\frac{k_{m}}{m}\right)=S_{\mu}(-t).$
We will also make use of the following interpolation over the interval $(0,1)$ of the finite free $S$-transform of the polynomials $p_{m}$:
$\mathcal{S}^{(m)}_{p_{m}}\colon(0,1)\to\mathbb{R},\quad\mathcal{S}^{(m)}_{p_{m}}(v)=S^{(m)}_{p_{m}}(-\frac{\lceil{mv}\rceil}{m}),\quad v\in\left(0,1-\frac{r_{m}}{m}\right),$
where we recall that $\lceil{\cdot}\rceil$ is the ceiling function and $r_{m}$ is the multiplicity of 0 in $p_{m}$ (not to be confused with r(m)=p(m)m)r(m)=p(m)-m). In this regard, we again emphasize that this interpolated S transform $\mathcal{S}^{(m)}_{p_{m}}$ is only defined on the interval $(0,1-r_{m}/m)$, while $T^{(m)}_{p_{m}}$ is defined on the whole interval $(0,1)$. The relation (19) becomes
$T^{(m)}_{p_{m}}(1-v)=\frac{1}{\mathcal{S}^{(m)}_{p_{m}}(v)},\quad v\in\left(0,1-\frac{r_{m}}{m}\right).$
@@
Let $(p_{m})_{m\geq 1}$ be a converging sequence of monic polynomials with positive roots and increasing degrees $m$. Let $\mu$ be the limiting measure :
$\mu\llbracket p_{m}\rrbracket\underset{m,+\infty}{\longrightarrow}\mu,$
and assume that $\mu\neq\delta_{0}$. Then,
$[\nabla^{(m)}](\mathcal{S}^{(m)}_{p_{m}})(v)=-\partial_{v}{S}_{\mu}(-v)+o(1),\quad m\rightarrow+\infty,$
locally uniformly on $v\in(0,1-\mu(\{0\})$.
Remark 5.2.2.
From the statement of Theorem 5.2.1, it is expected that the iterated discrete derivatives of the finite $T$-transform converge to the iterated derivatives of the $T$ transform. We will not prove this assertion, a possible strategy would be to use the same arguments as exposed in the proof of Theorem 5.2.1 in a recursive way.
Corollary 5.2.3.
Continuing the setting of Theorem 5.2.1, if $\mu$ is the limit measure of the $p_{m}$, then for any $v\in(\mu{(\{0\})},1)$,
$\nabla^{(m)}](T^{(m)}_{p_{m}})(v)=\partial_{v}T_{\mu}(v-1)+o(1)$
as $m$ goes to infinity.
@@
Given that for any $v\in(\mu(\{0\}),1)$:
$T_{p_{m}}^{(m)}(v)=\frac{1}{\mathcal{S}^{(m)}_{p_{m}}(1-v)},$
we infer
$\nabla^{(m)}T^{(m)}_{p_{m}}(v)$ $=-m\frac{\mathcal{S}^{(m)}_{p_{m}}(1-v-\frac{1}{m})-\mathcal{S}^{(m)}_{p_{m}}(1-v)}{\mathcal{S}^{(m)}_{p_{m}}(1-v-\frac{1}{m})\mathcal{S}^{(m)}_{p_{m}}(1-v)}$
$=\frac{\tilde{\nabla}^{(m)}\mathcal{S}^{(m)}_{p_{m}}(1-v)}{\mathcal{S}^{(m)}_{p_{m}}(1-v-\frac{1}{m})\mathcal{S}^{(m)}_{p_{m}}(1-v)}\rightarrow_{m,+\infty}-\partial_{v}{S}_{\mu}(v-1)/{S}_{\mu}(v-1)^{2}$
where we have use the local uniform convergence of $\mathcal{S}^{(m)}_{p_{m}}$ to $\mathcal{S}_{\mu}$ (as an application of the second Dini’s theorem, since $S_{\mu}$ is monotone non-increasing) and 5.2.1 for the left derivative operator $\tilde{\nabla}^{(m)}$. Since
$\partial_{v}{S}_{\mu}(v-1)/{S}_{\mu}(v-1)^{2}=\partial_{v}T_{\mu}(v-1),$
the corollary is proved. ∎
Remark 5.2.4.
Albeit the pointwise convergence of the finite free $T$ transform $T^{(m)}_{p_{m}}$ holds over the interval $(0,1)$ [arizmendi2024s, Theorem 9.1], the convergence of its discrete derivatives is stated only over the interval $(\mu(\{0\}),1)$. The behaviour of the discrete derivatives over $(0,\mu(\{0\}))$ remains unclear.
Proof of Theorem 5.2.1.
We closely follow the line of arguments presented in [arizmendi2024s] to prove Propositions 7.2, 7.3 and 7.8 based on the following representation of the free $S$ transform. Let $v\in(0,1)$, we denote by ${\rm Dil}_{v}(\mu)$ the dilation by $v$ of the probability measure $\mu$, in particular:
$\int x^{\ell}{\rm Dil}_{v}(\mu)(dx)=v^{\ell}\int x^{\ell}\mu(dx).$
Recall also the definition of fractional additive convolution power [shlyakhtenko2020fractional] in terms of the free cumulants. If $\mu$ is compactly supported, then $\mu^{\boxplus(1/v)}$ is the unique compactly supported probability measure whose free cumulants are the $(1/v)$-multiples of those of $\mu$. With these definitions [belinschi2008remarkable]:
$S_{\mu}(-v)=-G_{{\rm Dil}_{v}(\mu^{\boxplus({1}/{v})})}(0),\quad v\in(0,1).$ (20)
where $G$ denotes the Cauchy transform. Finally, recall the definition of the operators $\partial_{k|m}$ acting in monic polynomials with degree $m$:
$\partial^{k|m}p_{m}=\frac{1}{(m)_{m-k}}p^{(m-k)},\quad 0\leq k\leq m.$ (21)
@@
Case 1: We make the additional assumption that there exists $\varepsilon>0$ and $\eta>0$ such that all the roots of the polynomials $p_{m}$ are contained in $\varepsilon,\eta)$.
Under this assumption, the convergence $\mu\llbracket p_{m}\rrbracket\to\mu$ is equivalent to the convergence of all the moments of $\mu\llbracket p_{m}\rrbracket\to\mu$ toward those of $\mu$ and further to the convergence of the finite free cumulants of $\mu\llbracket p_{m}\rrbracket\to\mu$ toward those of $\mu$. This is out of the scope of the present article to give an account about finite free cumulants, we refere the reader to [arizmendi2018cumulants].
@@
We begin by relating the finite free $S$-transform to the Cauchy transform of the counting measure of the roots of derivatives of $p_{m}$, (see [arizmendi2024s], Lemma 6.11 2). Let $v\in(0,1)$.
$m(\mathcal{S}^{(m)}_{p_{m}}(v+\frac{1}{m})-\mathcal{S}_{p_{m}}^{(m)}(v))$ $=m(\mathcal{S}^{(m)}_{p_{m}}(-\frac{\lceil{mv}\rceil+1}{m})-\mathcal{S}^{(m)}_{p_{m}}(-\frac{\lceil{mv}\rceil}{m}))$
$=m(G_{\mu\llbracket\partial^{\lceil{mv}\rceil|}p_{m}\rrbracket}(0)-G_{\mu\llbracket\partial^{\lceil{mv}\rceil+1|m}p_{m}\rrbracket}(0))$
$=-G_{\nu^{(m)}_{p_{m}}(v)}(0),$
@@
where $\nu^{(m)}_{p_{m}}(v):=\mu\llbracket\partial^{\lceil{mv}\rceil|}p_{m}\rrbracket-\mu\llbracket\partial^{\lceil{mv}\rceil+1|m}p_{m}\rrbracket$. Observe that this measure has total variation equal to $2d$, it is therefore not uniformly bounded. Nevertheless, we will study the convergence of its Cauchy transform by using the moment method. By definition, the $\ell^{th}$ moment $M_{\ell}(\nu_{p_{m}}^{(m)}(v))$ of $\nu_{p_{m}}^{(m)}(v)$ is
$M_{\ell}(\nu_{p_{m}}^{(m)}(v))$ $=m(M_{\ell}(\partial^{\lceil{mv}\rceil+1|m}p_{m})-M_{\ell}(\partial^{\lceil{mv}\rceil|m}p_{m}))$ (22)
@@
Recall the following asymptotic expansion between the moments $M_{\ell}(p),\,\ell\geq 1$ and the finite free cumulants $\kappa^{(m)}_{\ell}(p),\ell\geq 1$ of a monic polynomial $p$ of degree $m$ (see [arizmendi2018cumulants]):
$M_{\ell}(p)=\sum_{\pi\in{\rm NC}(\ell)}\kappa^{(m)}_{\pi}(p)+\frac{2\ell}{m}\sum_{\begin{subarray}{c}r+s=\ell\\ \pi\in S_{{\rm NC}}(r,s)\end{subarray}}r^{-1}s^{-1}\kappa^{(m)}_{\pi}(p)+O(\frac{1}{m^{2}})$ (23)
@@
and the following simple formula relating the finite free cumulant of a polynomial $p$ and its derivatives $\partial^{k|m}p$ (see [arizmendi2024s]) :
$\kappa^{(k)}_{\ell}(\partial^{k|m}(p))=(\frac{k}{m})^{\ell-1}\kappa^{(m)}_{\ell}(p),\quad\ell\leq k.$ (24)
Inserting (23) in (22) and using further equation (24), we obtain:
$M_{\ell}(\nu^{(m)}_{p_{m}}(v))$ $=m\bigg(\sum_{\pi\in{\rm NC}(\ell)}\bigg[(\frac{\lceil{mv}\rceil+1}{m})^{\ell-|\pi|}\kappa^{(m)}_{\pi}(p_{m})-(\frac{\lceil{mv}\rceil}{m})^{\ell-|\pi|}\kappa_{\pi}^{(m)}(p_{m})\bigg]$
$\hskip 56.9055pt+2\frac{\ell}{\lceil{mv}\rceil+1}\sum_{\begin{subarray}{c}r+s=\ell\\ \pi\in S_{{\rm NC}}(r,s)\end{subarray}}r^{-1}s^{-1}(\frac{\lceil{mv}\rceil+1}{m})^{\ell-|\pi|}\kappa^{(m)}_{\pi}(p_{m})$
@@
$\hskip 142.26378pt-2\frac{\ell}{\lceil{mv}\rceil}\sum_{\begin{subarray}{c}r+s=\ell\\ \pi\in S_{{\rm NC}}(r,s)\end{subarray}}r^{-1}s^{-1}(\frac{\lceil{mv}\rceil}{m})^{\ell-|\pi|}\kappa_{\pi}^{(m)}(p_{m})\bigg)$
$+O(1/m)$
@@
For the first term in the right-hand side of the equation above, with $\kappa_{\pi}(\mu),\pi\in{\rm{\rm NC}}(\ell)$ the partitioned free cumulants of $\mu$:
$m\sum_{\pi\in{\rm NC}(\ell)}\bigg[(\frac{\lceil{mv}\rceil+1}{m})^{\ell-|\pi|}$ $\kappa^{(m)}_{\pi}(p)-(\frac{\lceil{mv}\rceil}{m})^{\ell-|\pi|}\kappa_{\pi}^{(m)}(p)\bigg]$
$=m\sum_{\pi\in{\rm NC}(\ell)}\bigg[(\frac{\lceil{mv}\rceil}{m})^{\ell-|\pi|}((1+\frac{1}{\lceil{mv}\rceil})^{\ell-|\pi|}-1)\kappa^{(m)}_{\pi}(p_{m})\bigg]$
$=\sum_{\begin{subarray}{c}\pi\in{\rm NC}(\ell)\\ \pi\neq\hat{0}_{\ell}\end{subarray}}\bigg[(\ell-|\pi|)v^{\ell-|\pi|-1}\kappa_{\pi}(\mu)\bigg]+O(1/m)$
@@
The sum of the second term and the third term are easily seen to contribute to a factor $O(\frac{1}{m})$. Hence, we obtain for the $\ell^{th}$ moment of $\nu_{p_{m}}^{(m)}(v)$ the following asymptotic expansion:
$M_{\ell}(\nu_{p_{m}}^{(m)}(v))$ $=\sum_{\begin{subarray}{c}\pi\in{\rm NC}(\ell)\\ \pi\neq\hat{0}_{\ell}\end{subarray}}\bigg[(\ell-|\pi|)v^{\ell-|\pi|-1}\kappa_{\pi}(\mu)\bigg]+O({1}/{m})$
$=\partial_{v}\sum_{\pi\in{\rm NC}(\ell)}v^{\ell-|\pi|}\kappa_{\pi}(\mu)+O(1/m)$
$=\partial_{v}m_{\ell}({\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}})+O(1/m)$
$=m_{\ell}(\partial_{v}{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}})+O(1/m)$
where in the last line $\partial_{v}{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}}$ is the weak limit of
$\frac{1}{\varepsilon}({\rm Dil}_{v+\varepsilon}\mu^{\boxplus\frac{1}{v+\varepsilon}}-{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}})$
as $\varepsilon$ goes to zero. Under our assumptions, weak convergence is equivalent to the convergence of the moments (as an easy application of the Ascoli’s theorem) and thus $\partial_{v}{\rm Dil}_{v}(\mu^{\boxplus 1/v})$ is the unique compactly supported signed measure such that
$M_{\ell}(\partial_{v}{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}})=\partial_{v}M_{\ell}({\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}})$
@@
Let us argue now that the Cauchy transform of $\nu_{p_{m}}^{(m)}(v)$ converges to the Cauchy transform of $\partial_{v}{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}}$ (analytical in the complementary of an interval $\varepsilon(v),\eta(v)]$ with $\varepsilon(v)>0$ (resp. $\eta(v)$) as a consequence of Lemma 5.6 and the proof of Proposition 5.5 in [arizmendi2024s]; in the sequel $\varepsilon$ (resp. $\eta$) will denote the minimum (resp. maximum) between $\varepsilon(v)$ and $\varepsilon>0$ (resp. between $\eta$ and $\eta(v)$)). We observe first that:
$G_{\nu_{p}(v)}=\{G_{\nu^{(m)}_{p_{m}}(v)},\,m\geq 1\}$
@@
is a normal family of holomorphic functions on the domain $\Omega=[\varepsilon,1]^{c}\subset\hat{\mathbb{C}}$$\mathbb{C}\cup\{$$\infty\}$, by applying Montel’s criterion (Fundamental Normality Test [schiff1993normal]). Let $G\colon\Omega\to\hat{\mathbb{C}}$ be an accumulation point of $G_{\nu_{p}(v)}$. First, $G$ is not identically equal to $\infty$ since $G_{\nu^{(m)}_{p_{m}}(v)}(\infty)=0$ for all $m\geq 1$. Since
$\frac{1}{\ell+1!}\frac{m^{\ell+1}}{dz^{\ell+1}}\biggr|_{z=0}G_{\nu^{(m)}_{p_{m}}(v)}(1/z)=m_{\ell}(\nu_{p_{m}^{(m)}})\rightarrow m_{\ell}(\partial_{v}{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}}),$